Please show all work and formulas for credit

Part 2 contains 3 Decision making recommendations that are independent of each other. Covers material in Module 17
Glass company manufactures glasses that it sells to mail-order distributors.
Sales price per pair of glasses: $62
Manufacturing and other costs follow:
Variable Cost per unit
Direct Materials $13
Direct labor 12
Factory overhead 2
Distribution 3
Total Variable costs $30
Fixed costs per month
Factory overhead $20,000
Selling and Administrative 10,000
Total Fixed costs $30,000
Current monthly production and sales volume 6,000 units
Monthly capacity 7000 units
Additional information: The variable distribution costs are for transportation to mail-order distributors.
Required: Determine the effect of each of the following 3 independent situations on monthly profits.
Make sure you show your computations to provide the monthly change in income and give a recommendation to management if they should accept the change.
#1 A $3 increase in the unit selling price should result in a 1,000-unit decrease in monthly sales
#2 A 10% decrease in the unit selling price should result in a 2,000-unit increase in monthly sales.
However, because of capacity constraints, the last 1,000 units would be produced during overtime with the
direct labor costs increasing by 40%.
#3 a British distributor has proposed to place a special one-time order for 1,000 units at a reduced
price of $57 per unit. The distributor would pay all transportation costs, so there are no variable distribution costs.  
There would be additional fixed selling and administrative costs of $1,000.
Answer #1

#1 We Should accept #1 plan we wiil get extra $2 per unit profit As per Below Satetment:-

Current # 1
Description Amoutn in $ Amoutn in $
Units Sales (A)         6,000.00         5,000.00
Sales Price Per unit (B)               62.00               65.00
Total Veriable Cost Per Unit (C)               30.00               30.00
Contribution Per Unit :- D= (B-C)               32.00               35.00
       
Fix cost Per Month ('E)      30,000.00      30,000.00
Net Profit Per Unit (F)
Total Contribution G= (D*A) 1,92,000.00 1,75,000.00
Total Profit for Month H= (G-E) 1,62,000.00 1,45,000.00
Profit Per Unit---------------> H/A 27 29

#2 We should not go for #2 plan because our profit per unit in decreasing by $ 10.29/- however units is sold increased as per below statement:-

Current # 2
Description Amoutn in $ Amoutn in $
Units Sales (A)         6,000.00         7,000.00
Sales Price Per unit (B)               62.00               55.80
Total Veriable Cost Per Unit (C)               30.00               30.00
Add Increase in Labour Cost 40%                 4.80
Contribution Per Unit :- D= (B-C)               32.00               21.00
       
Fix cost Per Month ('E)      30,000.00      30,000.00
Net Profit Per Unit (F)
Total Contribution G= (D*A) 1,92,000.00 1,47,000.00
Total Profit for Month H= (G-E) 1,62,000.00 1,17,000.00
Profit Per Unit---------------> H/A 27               16.71

#3 We should not go for plan #3 as due to this we will get extra loss of $ 4000 for 1000 units as per below statement:-

# 3
Description Amoutn in $
Units Sales (A)         1,000.00
Sales Price Per unit (B)               57.00
Total Veriable Cost Per Unit (C)               30.00
Add Increase in Labour Cost 40%                      -  
Contribution Per Unit :- D= (B-C)               27.00
   
Fix cost Per Month ('E) With Extra $ 1000      31,000.00
Net Profit Per Unit (F)
Total Contribution G= (D*A)      27,000.00
Total Profit for Month H= (G-E)       -4,000.00
Profit Per Unit---------------> H/A               -4.00